# Relationship Between Major Stellar Physical Parameters and Normal Mode Frequencies in Accreting White Dwarf Stars

### v1

@@ Abstract
White dwarfs (WDs) are the final fate of about 97% of the stars in our galaxy, making them vital tracers of stellar history. A fraction of WDs exist in cataclysmic variable (CV) systems, accreting matter from a nearby companion star. A subset of CVs undergo episodic rapid mass transfer, termed dwarf novae (DNe) outbursts. Some accreting WDs exhibit near sinusoidal photometric variations, interpreted as $g$ mode pulsations. However, identifying pulsation modes in accreting WDs remains challenging due to the paucity of available observed modes. In this work, we present a comprehensive computation of the observable $g$ mode frequencies across a range of WD parameters, varying the WD mass, size of the newly accreted layer and core temperature. We also introduce a novel method for mode identification based on the time evolution of pulsation periods following an accretion episode. Our mode identification method does not rely on the direct detection of the consecutive radial mode orders, frequently required in isolated WDs. Moreover, this work improves upon our previous WD modeling efforts. We use a more realistic core temperature in addition to thermohaline mixing and element diffusion enabled during the accretion phase.
@@ 1 Introduction
While pulsations in isolated WDs like ZZ Ceti and V777 Her have been extensively studied (Winget & Kepler, 2008; Fontaine & Brassard, 2008; Aerts et al., 2010; Althaus et al., 2010; Romero et al., 2012), our understanding of pulsations in rapidly rotating accreting WDs remains limited (Townsley et al., 2004; Saio, 2019; Kumar & Townsley, 2023, 2025). In this work, we aim to continue to lay the groundwork for interpreting the oscillations seen in accreting WDs through a forward modeling approach. WD asteroseismology holds a great potential to reveal intrinsic stellar properties by constraining the thickness of the newly accreted mixed layer, stellar mass, properties of the core composition layers, and more.
@@ 2 Method
In this section, we describe the making of the WD models and elaborate on the choice of parameters used in this study. Sections 2.1 to 2.3 extend the discussion of the evolutionary stages of the WD, along with highlighting differences from our previous works (Kumar & Townsley, 2023, 2025). Section 2.4 outlines the choice of parameter space used for the WD models and how it relates to what is known for the CV WDs. Section 2.5 outlines our choice of modes that we expect might be seen prominently on accreting WDs.
@@ 2.1 WD Models Before Accretion
Here we describe the detailed modeling of the CV WDs and the models used to compute the normal modes. We use the one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019), to perform our stellar evolutionary calculations. The general layout of constructing the WD is similar to what is used in Timmes et al. (2018); Kumar & Townsley (2023, 2025). In total, we consider five different WD mass models. Two of these are adopted from Kumar & Townsley (2023), with their final masses of 0.78 $M_{\odot}$ (initial mass = 3.95 $M_{\odot}$) and 0.93 $M_{\odot}$ (initial mass = 6.1 $M_{\odot}$). While we used their models after the Asymptotic Giant Phase (AGB) as initial configurations, we cooled them using the more extended pp_cno_extras_o18_ne22.net reaction network including the element diffusion, rather than the basic_network.net employed in their work. In addition, all models in this study were computed with MESA release version r15140, which provides improved timestep control compared to the version r10398 used by Kumar & Townsley (2023).
@@ 2.2 Long-Term Time Averaged Accretion
The freshly accreted hydrogen and helium in a mass-transferring binary system induce an unstable thermonuclear ignition on the surface, resulting in a Classical Nova (CN) outburst (Townsley & Bildsten, 2004; Nomoto et al., 2007; Schreiber et al., 2016). The cooled WD is allowed to settle via further cooling for $\sim 10^{5}$ years without accretion or element diffusion. Accretion is then initiated, still without element diffusion, for another $\sim 10^{5}$ years (relatively short period compared to overall WD lifetime) and with convection turned off. This prevent any premixing between the freshly accreted material and the helium remnants on the WD’s surface leftover from the previous evolutionary stage. The interface between the layers created by this process is not critically important anyway since it is thin enough to be smoothed out by later diffusion. Following this, the WD is subjected to long-term accretion with rate, $\dot{M}$, chosen to give the desired physical state as described in section 2.4, with element diffusion and predictive mixing in place during this accretion phase. Unless otherwise required by parameter choices, we use a rate of $\dot{M}=6\times 10^{-11}~M_{\odot}$ yr-1 or the rate appropriate to the relevant long-term $\dot{M}$ for that case. This approach is essential for ensuring a physically valid core-shell boundary structure, which directly influences mode frequency spacings. We perform this long-term accretion on all the five different WD mass models, along with additional lower $T_{\rm c}$ models of a 0.78 $M_{\odot}$ (see below for details on our parameter choices). Rotation is included from the beginning through to the cooling phase of the WD. However, we do not include rotation during the accretion phase, instead, we explicitly incorporate rotational effects in GYRE (Townsend & Teitler, 2013; Townsend et al., 2018) when computing the normal mode frequencies.
@@
The freshly accreted layer, $M_{\rm acc}$, is parameterized by the mass. The accreted mass fractions are $X_{\rm H}=0.742$, $X_{\mathrm{{4_{He}}}}=0.237$, $X_{\rm Z}=0.02$, and $X_{\mathrm{{3_{He}}}}=0.003$. As described by Townsley & Bildsten (2004), a trace of $^{3}{\rm He}$ from the donor (Shara, 1980) can lead to early ignition at low temperatures. We use the mixing length parameter $\alpha_{\mathrm{mlt}}=1.5$, $\alpha_{\mathrm{semi-convection}}=0.01$, and $\alpha_{\mathrm{thermohaline}}=1$. Due to persistent numerical challenges in modeling multiple hydrogen shell flashes with convective boundary mixing, we instead focus on accretion onto a bare white dwarf and evaluate its seismological properties prior to the onset of the hydrogen shell flash. This alternative is preferable to simulating an unrealistic, numerically unstable, and unconstrained mixed layer. Therefore, we halt the accretion evolution when the hydrogen surface luminosity of the WD $\log L_{\rm H}\simeq$ 5 $L_{\odot}$ (where $L_{\odot}$ is the solar luminosity), accumulating mass immediately prior to the explosion or the next CN cycle, $M_{\rm ign}$. The maximum envelope mass $M_{\rm acc}$ leading to $M_{\rm ign}$ is set by the unstable nuclear burning. This process is mostly governed by the proton-proton (ppp-p) chain, near the base of the accreted H/He layer that causes the CN runaway. As the accreted material builds up on the surface, the temperature and pressure at the base of the envelope increase, finally triggering the runaway, marking the ignition point with setting the upper limit on $M_{\rm acc}$ before the outburst. To make a note, we do not use the ignition model, it is only used to set the $M_{\rm ign}$ for making models a comparable fraction toward ignition.
@@ 2.3 Short-Term Accretion or Dwarf Novae
Dwarf novae (DNe) are the subset of the CVs (Warner, 1995) that undergo dramatic, episodic accretion events onto the WD surface. These outbursts last from a few days to months timescales, with a high accretion rate of $\sim 10^{-8}M_{\odot}$ yr-1, followed by a quiescent state lasting from days to decades. During quiescence, the accretion rate onto the WD drops significantly, allowing a more direct inspection of the WD surface, most notably $T_{\rm eff}$. A far-ultraviolet flux spectroscopic examination of the underlying accreting WDs reveals that they cool to their pre-outburst quiescent temperatures (Pringle, 1988; Long et al., 1993; Szkody et al., 2012), a phenomenon that has also been supported by theoretical models (Piro et al., 2005; Kumar & Townsley, 2023).
@@
Figure 1: Temperature profiles against the fractional depth ($\log(1-m/M)$) at the beginning of the long-term accretion phase (top panel, solid) and the hydrogen luminosity $\log(L_{\rm H}/L_{\odot})=2$ (top panel, dashed)and 5 (bottom panel, dot-dashed), for three different core temperatures with $T_{\rm c}=5,7$, and $9\times 10^{6}$ K of a 0.78 $M_{\odot}$ WD model.
Figure 2: WD compositions and frequency profiles after a long-term accretion phase with $M_{\text{acc}}=1.5\times 10^{-4}~M_{\odot}$ at core temperatures $T_{\rm c}=5$ and 7 $\times 10^{6}$ K of a 0.78 $M_{\odot}$ WD model at $T_{\rm eff}=14000$ K. Brunt-Väisälä and Lamb frequency are shown in the bottom panel. Mass fractions are shown on the log scale in the middle panel to highlight the other significant species produced that are difficult to note on the linear scale (top panel). The fraction of solid core is denoted by the vertical dashed line and is about 34% in radius for $T_{\rm c}=5\times 10^{6}$ K. The WD is entirely liquid at $T_{\rm c}=7\times 10^{6}$ K.
@@ 2.4 Description of the Parameter Space
We now present a comprehensive description of the parameter space employed in this study, along with the rationale behind our choices. Thanks to the $Gaia$ mission (Gaia Collaboration et al., 2016, 2018) and the ultraviolet data from the HubbleSpaceTelescopeHubble~Space~Telescope, the number of CVs observed with accurately determined WD masses, $M_{\rm WD}$ and $T_{\rm eff}$, has significantly increased, now exceeding over one hundred (Pala et al., 2017; Shara et al., 2018; Pala et al., 2022). CV binaries are believed to lose their angular momentum through either gravitational radiation or magnetic breaking depending on their orbital period (Kolb & Baraffe, 1999; Howell et al., 2001; Townsley & Bildsten, 2003). This angular momentum loss determines the average accretion rate $\langle\dot{M}\rangle$ onto the WD. The WD’s $T_{\rm eff}$ traces this process, as it reflects compressional heating from the accreted material (Townsley & Bildsten, 2004; Townsley & Gänsicke, 2009), thereby constraining $\langle\dot{M}\rangle$ (Townsley & Bildsten, 2003). A relatively good estimates of $\langle\dot{M}\rangle$ may inform the accumulated mass, $M_{\rm acc}$, and the ignition mass, $M_{\rm ign}$(Townsley & Bildsten, 2004). Hence, we anchor our study on this crucial parameter, $T_{\rm eff}$.
We break our seismological study into three independent scenarios and with: (1) Models with diverse core temperatures, $T_{\rm c}$, while keeping the accreted layer mass, $M_{\rm acc}$, fixed for a 0.78 $M_{\odot}$ mass model, (2) varying $M_{\rm acc}$ at a fixed $T_{\rm c}=5\times 10^{6}$ K for a 0.78 $M_{\odot}$ mass model, and (3) models with five different WD masses, $M_{\rm WD}$ at a fixed $T_{\rm c}=6\times 10^{6}$ K. In each individual case, $\langle\dot{M}\rangle$ is adjusted to give a $T_{\rm eff}$ of 14000 K. In our previous work, (Kumar & Townsley, 2023), we adopted a lower accretion rate ($\sim 10^{-12}M_{\odot}$ yr-1) and correspondingly higher $T_{\rm c}$ ($\sim 10^{7}$ K). However, the recent results from Pala et al. (2022) suggests that the WDs in CVs below the period gap accretes at rates higher than previously assumed. Based on their reported $\langle\dot{M}\rangle$ and $\langle\dot{M}\rangle$ derived from $T_{\rm eff}$ (Townsley & Bildsten, 2003), WDs in CVs are expected to have cooler $T_{c}$ (Townsley & Bildsten, 2004). Therefore, in this study, we adopt an accretion rate of $\sim 10^{-10}M_{\odot}$ yr-1, consistent with expectations for CVs at their current evolutionary stage (Townsley & Gänsicke, 2009), correspondingly lower $T_{\rm c}$ in order to match observed $T_{\rm eff}$ (Townsley & Bildsten, 2004). For a 0.78 $M_{\odot}$, the equilibrium core temperatures, $T_{\text{c, eq}}$, falls within a range of $5-7\times 10^{6}$ K.
@@ 2.4.1 Varying Central Temperature
The model parameters are summarized in Table 1. It is important to note that we do not actively investigate the $T_{\text{c, eq}}$ for a given value of $\langle\dot{M}\rangle$ by evolving through a large number of H shell flashes (See discussions in Townsley & Bildsten (2004)). However, once the WD core equilibrate, the observable quantities remain largely insensitive to $M_{\rm WD}$ and $\langle\dot{M}\rangle$. Figure 1 presents the temperature profiles of a 0.78 $M_{\odot}$ WD as a function of fractional depth ($\log(1-m/M)$). The profiles are shown for three core temperatures, $T_{\rm c}=5,7$ and $9\times 10^{6}$ K represented with red, blue, and green lines, respectively. The top panel shows the temperature profiles at the beginning of the long-term accretion phase (solid lines) and when the surface hydrogen luminosity reaches $L_{\rm H}=10^{2}~L_{\odot}$ (dashed lines). The bottom panel display the profiles at slightly advanced phase of the WD evolution (well past the shell’s ignition point), when $L_{\rm H}=10^{5}~L_{\odot}$ (dash-dotted lines). We see from the interior most part of the temperature profile shown in Figure 1 that the WD center will experience significant cooling for the $T_{\rm c}=9\times 10^{6}$ K model during the accumulation of the surface layer and remains in overall net cooling phase even as ignition approaches. However, at lower $T_{\rm c}=5\times 10^{6}$ K, after cooling initially, the temperature profile changes enough to stop the cooling of or possibly slightly reheat the center as $M_{\rm acc}$ increases. This feature suggests that the WD equilibrium core temperature should lie between 5 and $7\times 10^{7}$ K, where core cooling is balanced by the heating for a given $\langle\dot{M}\rangle$ across the CN cycle. However, running a full CN evolutionary model with both overshooting and element diffusion enabled in MESA does not appear to be currently feasible, as they produce nonphysical profiles due to the various mixing processes in place. This numerical artifact is particularly evident in the bottom panel of Figure 1 for the $T_{\rm c}=5\times 10^{6}$ K. Due to these limitations, we are currently unable to perform a full CN cycle that includes both overshooting and element diffusion.
@@
We construct five different WD models with $T_{\text{c}}=5,5.5,6,6.5,\text{and}~7\times 10^{6}$ K for a 0.78 $M_{\odot}$ WD. In order to determine the approximate ignition mass, the different $T_{c}$ models are subjected to long-term accretion with $\langle\dot{M}\rangle=6\times 10^{-11}M_{\odot}$ yr-1, resulting in ignition masses of $M_{\mathrm{ign}}=3.57,3.05,2.84,2.05,$ and $1.69\times 10^{-4}M_{\odot}$ for $T_{\rm c}=5,5.5,6,6.5,$ and $7\times 10^{6}$ K, respectively for the 0.78 $M_{\odot}$ model. To ensure consistency, we then select models with a uniform accreted mass $M_{\rm acc}=1.5\times 10^{-4}M_{\odot}$ for all $T_{\rm c}$ values. At this $M_{\rm acc}$, we determine an appropriate $\langle\dot{M}\rangle$ that yields the star surface temperature of $T_{\rm eff}=14000$ K, for all $T_{\rm c}$ values. We find $\langle\dot{M}\rangle=12.37,11.42,11.10,9.42,$ and $7.94\times 10^{-11}M_{\odot}~{\rm yr}^{-1}$, respectively for $T_{\rm c}=5,5.5,6,6.5,$ and $7\times 10^{6}$ K. Some of these $T_{\rm c}$ are low enough that the plasma Coulomb parameter $\Gamma$ exceeds the threshold for forming a solid ($\Gamma\gtrsim 175$, Isern et al., 1997; Potekhin & Chabrier, 2000, 2010). Consequently, a fraction of the WD core solidifies, with the extent of crystallization depending on both $T_{\rm c}$ and the WD mass, $M_{\rm WD}$. For a 0.78 $M_{\odot}$ WD, the solid core fraction contributes approximately 26 (34), 12 (25), and 3 (15)% cores that are solid in mass (radius) coordinates, for $T_{\rm c}=5,5.5,$ and $6\times 10^{6}$ K. However, at higher core temperatures $T_{\rm c}=6.5$ and $5.5\times 10^{6}$ K, the WD core remains entirely liquid.
@@
Figure 2 illustrates the WD thermal structure following the long-term accretion phase with $M_{\rm acc}=1.5\times 10^{-4}M_{\odot}$ for $T_{\rm c}=5~(solid~lines)$ and $7\times 10^{6}~(dashed~lines)$ K for a 0.78 $M_{\odot}$ at $T_{\rm eff}=14000$ K. The top two panels display the mass fractions of the key elemental species, with the middle panel shown on a logarithmic scale to display the significant contributions from the other less abundant but important species. The vertical dashed line (black) indicates the $\log P$ at which the core is completely solid for $T_{\rm c}=5\times 10^{6}$ K. The bottom panel presents the buoyancy and lamb frequency profiles. The peak in buoyancy frequency at $\log P\approx 19$ $\mathrm{erg~cm^{-3}}$ reflects the transition between the hydrogen-rich layer and helium. Another peak located between $\log P$ of $\sim$ 23 and 24 is contributed due to the CO gradient at the edge of the remnant of the convective core formed during core He burning. Notably, the overall neon abundances remain the same irrespective of WD core temperatures. The slight discontinuity at $\log P$ of $\sim 15.8$ corresponds to the outermost layer in which thermohaline mixing is effective due to the finite timestep in the simulation. While the presence or absence of a solid core does not lead us to any obvious differences in the internal structural features, asteroseismology may be able to provide insights into the extent of core crystallization through the forward modeling method, as we continue to study here.

### v2

@@ Abstract
White dwarfs (WDs) are the final fate of about 97% of the stars in our galaxy, making them vital tracers of stellar history. A fraction of WDs exist in cataclysmic variable (CV) systems, accreting matter from a nearby companion star. A subset of CVs undergo episodic rapid mass transfer, termed dwarf novae (DNe) outbursts. Some accreting WDs exhibit near sinusoidal photometric variations, interpreted as $g$-mode pulsations. However, identifying pulsation modes in accreting WDs remains challenging due to the paucity of available observed modes. In this work, we present a comprehensive computation of the observable $g$-mode frequencies across a range of WD parameters, varying the WD mass, size of the newly accreted layer and core temperature. We also introduce a novel method for mode identification based on the time evolution of pulsation periods following an accretion episode. Our mode identification method does not rely on the direct detection of the consecutive radial mode orders, frequently required in isolated WDs. Moreover, this work improves upon our previous WD modeling efforts. We use a more realistic core temperature in addition to thermohaline mixing and element diffusion enabled during the accretion phase.
@@ 1 Introduction
While pulsations in isolated WDs like ZZ Ceti and V777 Her have been extensively studied (Winget & Kepler, 2008; Fontaine & Brassard, 2008; Aerts et al., 2010; Althaus et al., 2010; Romero et al., 2012), our understanding of pulsations in rapidly rotating accreting WDs remains limited (Townsley et al., 2004; Saio, 2019; Kumar & Townsley, 2023, 2026). In this work, we aim to continue to lay the groundwork for interpreting the oscillations seen in accreting WDs through a forward modeling approach. WD asteroseismology holds a great potential to reveal intrinsic stellar properties by constraining the thickness of the newly accreted mixed layer, stellar mass, properties of the core composition layers, and more.
@@ 2 Method
In this section, we describe the making of the WD models and elaborate on the choice of parameters used in this study. Sections 2.1 to 2.3 extend the discussion of the evolutionary stages of the WD, along with highlighting differences from our previous works (Kumar & Townsley, 2023, 2026). Section 2.4 outlines the choice of parameter space used for the WD models and how it relates to what is known for the CV WDs. Section 2.5 outlines our choice of modes that we expect might be seen prominently on accreting WDs.
@@ 2.1 WD Models Before Accretion
Here we describe the detailed modeling of the CV WDs and the models used to compute the normal modes. We use the one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019), to perform our stellar evolutionary calculations. The general layout of constructing the WD is similar to what is used in Timmes et al. (2018); Kumar & Townsley (2023, 2026). In total, we consider five different WD mass models. Two of these are adopted from Kumar & Townsley (2023), with their final masses of 0.78 $M_{\odot}$ (initial mass = 3.95 $M_{\odot}$) and 0.93 $M_{\odot}$ (initial mass = 6.1 $M_{\odot}$). While we used their models after the Asymptotic Giant Phase (AGB) as initial configurations, we cooled them using the more extended pp_cno_extras_o18_ne22.net reaction network including the element diffusion, rather than the basic_network.net employed in their work. In addition, all models in this study were computed with MESA release version r15140, which provides improved timestep control compared to the version r10398 used by Kumar & Townsley (2023).
@@ 2.2 Long-Term Time Averaged Accretion
The freshly accreted hydrogen and helium in a mass-transferring binary system induce an unstable thermonuclear ignition on the surface, resulting in a Classical Nova (CN) outburst (Townsley & Bildsten, 2004; Nomoto et al., 2007; Schreiber et al., 2016). The cooled WD is allowed to settle via further cooling for $\sim 10^{5}$ years without accretion or element diffusion. Accretion is then initiated, still without element diffusion, for another $\sim 10^{5}$ years (relatively short period compared to overall WD lifetime) and with convection turned off. This prevent any premixing between the freshly accreted material and the helium remnants on the WD’s surface leftover from the previous evolutionary stage. The interface between the layers created by this process is not critically important anyway since it is thin enough to be smoothed out by later diffusion. Following this, the WD is subjected to long-term accretion with rate, $\dot{M}$, chosen to give the desired physical state as described in section 2.4. Element diffusion and thermohaline mixing are included throughout the accretion phase. We also enable predictive mixing during accretion, as it helps stabilize the intermittent mixing regions that result from the interplay of accretion, diffusion, and thermohaline. Unless otherwise required by parameter choices, we use a rate of $\dot{M}=6\times 10^{-11}~M_{\odot}$ yr-1 or the rate appropriate to the relevant long-term $\dot{M}$ for that case. This approach is essential for ensuring a physically valid core-shell boundary structure, which directly influences mode frequency spacings. We perform this long-term accretion on all the five different WD mass models, along with additional lower $T_{\rm c}$ models of a 0.78 $M_{\odot}$ (see below for details on our parameter choices). Rotation is included from the beginning through to the cooling phase of the WD. However, we do not include rotation during the accretion phase, instead, we explicitly incorporate rotational effects in GYRE (Townsend & Teitler, 2013; Townsend et al., 2018) when computing the normal mode frequencies.
@@
The freshly accreted layer, $M_{\rm acc}$, is parameterized by the mass. The accreted mass fractions are $X_{\rm H}=0.742$, $X_{\mathrm{{4_{He}}}}=0.237$, $X_{\rm Z}=0.02$, and $X_{\mathrm{{3_{He}}}}=0.003$. As described by Townsley & Bildsten (2004), a trace of $^{3}{\rm He}$ from the donor (Shara, 1980) can lead to early ignition at low temperatures. We use the mixing length parameter $\alpha_{\mathrm{mlt}}=1.5$, $\alpha_{\mathrm{semi-convection}}=0.01$, and $\alpha_{\mathrm{thermohaline}}=1$. Due to persistent numerical challenges in modeling multiple hydrogen shell flashes with convective boundary mixing, we instead focus on accretion onto a bare white dwarf and evaluate its seismological properties prior to the onset of the hydrogen shell flash. This alternative is preferable to simulating an unrealistic, numerically unstable, and unconstrained mixed layer. Therefore, we halt the accretion evolution when the hydrogen surface luminosity of the WD $\log L_{\rm H}\simeq$ 5 $L_{\odot}$ (where $L_{\odot}$ is the solar luminosity), accumulating mass immediately prior to the explosion or the next CN cycle, $M_{\rm ign}$. The maximum envelope mass $M_{\rm acc}$ leading to $M_{\rm ign}$ is set by the unstable nuclear burning. This process is mostly governed by the proton-proton (ppp-p) chain, near the base of the accreted H/He layer that causes the CN runaway. As the accreted material builds up on the surface, the temperature and pressure at the base of the envelope increase, finally triggering the runaway, marking the ignition point with setting the upper limit on $M_{\rm acc}$ before the outburst. To make a note, we do not use the ignition model; it is only used to set the $M_{\rm ign}$, so that the models correspond to a comparable fraction of the path toward ignition.
@@ 2.3 Short-Term Accretion or Dwarf Novae
Dwarf novae (DNe) are the subset of the CVs (Warner, 1995) that undergo dramatic, episodic accretion events onto the WD surface. These outbursts last from a few days to months, with a high accretion rate of $\sim 10^{-8}M_{\odot}$ yr-1, followed by a quiescent state lasting from days to decades. During quiescence, the accretion rate onto the WD drops significantly, allowing a more direct inspection of the WD surface, most notably $T_{\rm eff}$. A far-ultraviolet flux spectroscopic examination of the underlying accreting WDs reveals that they cool to their pre-outburst quiescent temperatures (Pringle, 1988; Long et al., 1993; Szkody et al., 2012), a phenomenon that has also been supported by theoretical models (Piro et al., 2005; Kumar & Townsley, 2023).
@@
Figure 1: Temperature profiles against the fractional depth ($\log(1-m/M)$) at the beginning of the long-term accretion phase (top panel, solid) and the hydrogen luminosity $\log(L_{\rm H}/L_{\odot})=2$ (top panel, dashed), pre-outburst (bottom panel, solid), and $\log(L_{\rm H}/L_{\odot})=5$ (bottom panel, dotted), for three different core temperatures with $T_{\rm c}=5,7$, and $9\times 10^{6}$ K of a 0.78 $M_{\odot}$ WD model.
Figure 2: WD compositions and frequency profiles after a long-term accretion phase with $M_{\text{acc}}=1.5\times 10^{-4}~M_{\odot}$ at core temperatures $T_{\rm c}=5$ and 7 $\times 10^{6}$ K of a 0.78 $M_{\odot}$ WD model at $T_{\rm eff}=14000$ K. Brunt-Väisälä and Lamb frequency are shown in the bottom panel. Mass fractions are shown on the log scale in the middle panel to highlight the other significant species produced that are difficult to note on the linear scale (top panel). The fraction of solid core is denoted by the vertical dashed line and is about 34% in radius for $T_{\rm c}=5\times 10^{6}$ K. The WD is uncrystallized at $T_{\rm c}=7\times 10^{6}$ K.
@@ 2.4 Description of the Parameter Space
We now present a comprehensive description of the parameter space employed in this study, along with the rationale behind our choices. Thanks to the $Gaia$ mission (Gaia Collaboration et al., 2016, 2018) and the ultraviolet data from the HubbleSpaceTelescopeHubble~Space~Telescope, the number of CVs observed with accurately determined WD masses, $M_{\rm WD}$ and $T_{\rm eff}$, has significantly increased, now exceeding over one hundred (Pala et al., 2017; Shara et al., 2018; Pala et al., 2022). CV binaries are believed to lose their angular momentum through either gravitational radiation or magnetic breaking depending on their orbital periods (Kolb & Baraffe, 1999; Howell et al., 2001; Townsley & Bildsten, 2003). This angular momentum loss determines the average accretion rate $\langle\dot{M}\rangle$ onto the WD. The WD’s $T_{\rm eff}$ traces this process, as it reflects compressional heating from the accreted material (Townsley & Bildsten, 2004; Townsley & Gänsicke, 2009), thereby constraining $\langle\dot{M}\rangle$ (Townsley & Bildsten, 2003). A relatively good estimates of $\langle\dot{M}\rangle$ may inform the accumulated mass, $M_{\rm acc}$, and the ignition mass, $M_{\rm ign}$(Townsley & Bildsten, 2004). Hence, we anchor our study on this crucial parameter, $T_{\rm eff}$.
We break our seismological study into three independent scenarios with: (1) Models with diverse core temperatures, $T_{\rm c}$, while keeping the accreted layer mass, $M_{\rm acc}$, fixed for a 0.78 $M_{\odot}$ mass model, (2) varying $M_{\rm acc}$ at a fixed $T_{\rm c}=5\times 10^{6}$ K for a 0.78 $M_{\odot}$ mass model, and (3) models with five different WD masses, $M_{\rm WD}$ at a fixed $T_{\rm c}=6\times 10^{6}$ K. In each individual case, $\langle\dot{M}\rangle$ is adjusted to give a $T_{\rm eff}$ of 14000 K. In our previous work, (Kumar & Townsley, 2023), we adopted a lower accretion rate ($\sim 10^{-12}M_{\odot}$ yr-1) and correspondingly higher $T_{\rm c}$ ($\sim 10^{7}$ K). However, the recent results from Pala et al. (2022) suggests that the WDs in CVs below the period gap accrete at rates higher than previously assumed. Based on their reported $\langle\dot{M}\rangle$ and $\langle\dot{M}\rangle$ derived from $T_{\rm eff}$ (Townsley & Bildsten, 2003), WDs in CVs are expected to have cooler $T_{c}$ (Townsley & Bildsten, 2004). Therefore, in this study, we adopt an accretion rate of $\sim 10^{-10}M_{\odot}$ yr-1, consistent with expectations for CVs at their current evolutionary stage (Townsley & Gänsicke, 2009), correspondingly lower $T_{\rm c}$ in order to match observed $T_{\rm eff}$ (Townsley & Bildsten, 2004). For a 0.78 $M_{\odot}$, the equilibrium core temperatures, $T_{\text{c, eq}}$, falls within a range of $5-7\times 10^{6}$ K.
@@ 2.4.1 Varying Central Temperature
The model parameters are summarized in Table 1. It is important to note that we do not actively investigate the $T_{\text{c, eq}}$ for a given value of $\langle\dot{M}\rangle$ by evolving through a large number of H shell flashes (See discussions in Townsley & Bildsten (2004)). However, once the WD core equilibrate, the observable quantities remain largely insensitive to $M_{\rm WD}$ and $\langle\dot{M}\rangle$. Figure 1 presents the temperature profiles of a 0.78 $M_{\odot}$ WD as a function of fractional depth ($\log(1-m/M)$). The profiles are shown for three core temperatures, $T_{\rm c}=5,7$ and $9\times 10^{6}$ K represented with red, blue, and green lines, respectively. The top panel shows the temperature profiles at the beginning of the long-term accretion phase (solid lines) and when the surface hydrogen luminosity reaches $L_{\rm H}=10^{2}~L_{\odot}$ (dashed lines). The bottom panel display the profiles just before shell’s ignition or pre-outburst (solid lines, onset of strong convection) and slightly advanced phase of the WD evolution (well past the shell’s ignition point), when $L_{\rm H}=10^{5}~L_{\odot}$ (dotted lines). From the innermost region of temperature profile shown in Figure 1, we see that the WD center experience significant cooling for the $T_{\rm c}=9\times 10^{6}$ K model during the accumulation of the surface layer and remains in overall net cooling phase even as ignition approaches. In this case, the WD exhibits an outward temperature gradient between $\log(1-m/M)=-2$ and -1. However, for the lower temperature model, $T_{\rm c}=5\times 10^{6}$ K, the WD initially cools but subsequently the temperature profile changes enough to stop the cooling of or possibly slightly reheat the center as $M_{\rm acc}$ increases. For this model, the temperature gradient becomes inward over the same fractional mass depth of -2 to -1. This feature suggests that the WD equilibrium core temperature should lie between 5 and $7\times 10^{7}$ K, where core cooling is balanced by the heating for a given $\langle\dot{M}\rangle$ across the CN cycle. However, running a full CN evolutionary model with both overshooting and element diffusion enabled in MESA does not appear to be currently feasible, as they produce nonphysical profiles due to the various mixing processes in place. This numerical artifact is particularly evident in the bottom panel of Figure 1 for the $T_{\rm c}=5\times 10^{6}$ K. Due to these limitations, we are currently unable to perform a full CN cycle that includes both overshooting and element diffusion.
@@
We construct five different non-magnetic CV WD models with $T_{\text{c}}=5,5.5,6,6.5,\text{and}~7\times 10^{6}$ K for a 0.78 $M_{\odot}$ WD. In order to determine the approximate ignition mass, the different $T_{c}$ models are subjected to long-term accretion with $\langle\dot{M}\rangle=6\times 10^{-11}M_{\odot}$ yr-1, resulting in ignition masses of $M_{\mathrm{ign}}=3.57,3.05,2.84,2.05,$ and $1.69\times 10^{-4}M_{\odot}$ for $T_{\rm c}=5,5.5,6,6.5,$ and $7\times 10^{6}$ K, respectively for the 0.78 $M_{\odot}$ model. To ensure consistency, we select models with the same accreted mass $M_{\rm acc}=1.5\times 10^{-4}M_{\odot}$ for all $T_{\rm c}$ values, as this mass remains below the $M_{\rm ign}$ in every case. At this $M_{\rm acc}$, we determine an appropriate $\langle\dot{M}\rangle$ that yields the star surface temperature of $T_{\rm eff}=14000$ K, for all $T_{\rm c}$ values. We find $\langle\dot{M}\rangle=12.37,11.42,11.10,9.42,$ and $7.94\times 10^{-11}M_{\odot}~{\rm yr}^{-1}$, respectively for $T_{\rm c}=5,5.5,6,6.5,$ and $7\times 10^{6}$ K. Some of these $T_{\rm c}$ are low enough that the plasma Coulomb parameter $\Gamma$ exceeds the threshold for forming a solid ($\Gamma\gtrsim 175$, Isern et al., 1997; Potekhin & Chabrier, 2000, 2010). Consequently, a fraction of the WD core solidifies, with the extent of crystallization depending on both $T_{\rm c}$ and the WD mass, $M_{\rm WD}$. For a 0.78 $M_{\odot}$ WD, the solid core fraction contributes approximately 26 (34), 12 (25), and 3 (15)% cores that are solid in mass (radius) coordinates, for $T_{\rm c}=5,5.5,$ and $6\times 10^{6}$ K. However, at higher core temperatures $T_{\rm c}=6.5$ and $5.5\times 10^{6}$ K, the WD core remains uncrystallized.
@@
Figure 2 illustrates the WD thermal structure following the long-term accretion phase with $M_{\rm acc}=1.5\times 10^{-4}M_{\odot}$ for $T_{\rm c}=5~(solid~lines)$ and $7\times 10^{6}~(dashed~lines)$ K for a 0.78 $M_{\odot}$ at $T_{\rm eff}=14000$ K. The sloped abundance profiles in the H-rich layer reflect the combined effects of thermohaline mixing and element diffusion operating throughout the accretion phase. The top two panels display the mass fractions of the key elemental species, with the middle panel shown on a logarithmic scale to display the significant contributions from the other less abundant but important species. The vertical dashed line (black) indicates the $\log P$ at which the core is completely solid for $T_{\rm c}=5\times 10^{6}$ K. The bottom panel presents the buoyancy and lamb frequency profiles. The peak in buoyancy frequency at $\log P\approx 19$ $\mathrm{erg~cm^{-3}}$ reflects the transition between the hydrogen-rich layer and helium. Another peak located between $\log P$ of $\sim$ 23 and 24 is contributed due to the CO gradient at the edge of the remnant of the convective core formed during core He burning. Notably, the overall neon abundances remain the same irrespective of WD core temperatures. The slight discontinuity at $\log P$ of $\sim 15.8$ corresponds to the outermost layer in which thermohaline mixing is effective due to the finite timestep in the simulation. While the presence or absence of a solid core does not lead us to any obvious differences in the internal structural features, asteroseismology may be able to provide insights into the extent of core crystallization through the forward modeling method, as we continue to study here.